Dome Metals has credit sales of $378,000 yearly with credit terms of net 60 days, which is also the average collection period. Dome does not offer a discount for early payment, so its customers take the full 60 days to pay

. a. What is the average receivables balance? (Use a 360-day year.) AVERAGE RECIEBLES BALANCE

b. What is the receivables turnover? ? (Use a 360-day year.) RECIEVABLES TURNOVER

Dome Metals has credit sales of $126,000 yearly. If Dome offers a 4 percent discount for payment in 18 days, what would the average accounts receivable balance be? Assume all customers would pay on the last day of the discount period. (Use a 360-day year.)

AVERAE RECIEIVABLE BALANCE

Dome Metals has credit sales of $486,000 yearly with credit terms of net 90 days, which is also the average collection period.

a. Assume the firm offers a 4 percent discount for payment in 15 days and every customer takes advantage of the discount. Also assume the firm uses the cash generated from its reduced receivables to reduce its bank loans which cost 12 percent. What will the net gain or loss be to the firm if this discount is offered? (Use a 360-day year.)

NET CHANGE IN INCOME

b. Should the firm offer the discount?

Yes No Dome Metals has credit sales of $270,000 yearly with credit terms of net 90 days, which is also the average collection period. Assume the firm adopts new credit terms of 2/15, net 90 and all customers pay on the last day of the discount period. Any reduction in accounts receivable will be used to reduce the firm's bank loan which costs 12 percent. The new credit terms will increase sales by 20% because the 2% discount will make the firm's price competitive.

a. If Dome earns 15 percent on sales before discounts, what will be the net change in income if the new credit terms are adopted? (Use a 360-day year.) NET CHANGE IN INCOME

b. Should the firm offer the discount? No Yes
Answer #1

(1) a) calculation of avg. recivable balance when credit term is 60 days

Since in the question it is mentioned the average collection period is 60 days. So formula will be

(Net credit sales) x (Avg collection period / Number of days in year ) = (378,000) x (60/360 ) = $63,000

b) receivable turnover = (Net credit sales / Avg. recivable) = $378,000/ $63,000 = 6

................................................................................................................................................................................

2) a) what average account receivable be if discount for payment within 18 days is given and it is assumed all pay on 18th day . So 18 days is also avg collection period, so formula will be

(Net credit sales) x (Avg collection period / Number of days in year ) = (126,000) x (18 / 360 ) = $ 6300

...................................................................................................................................................................................

3) a) Dome Metals has credit sales of $486,000 yearly with credit terms of net 90 days, which is also the average collection period.

So formula will of avg. recivable balance will be

(Net credit sales) x (Avg collection period / Number of days in year ) = (486,000) x (90/360 ) = $121,500

when customer avail benifit of 4% discount of payment in 15 days

avg. recivable balance will be

(Net credit sales) x (Avg collection period / Number of days in year ) = (486,000) x (15/360 ) = $20,250

Excess cash company will have due to customer early payment = $121,500 - $ 20,250 = 101,250

The discount amount = 121,500x 4%= 4,860

Net excess cash company have due to customer early payment = 101,250 - 4,860 = $ 96,390

Bank Interst amount saved by company = $ 96,390 x 12% = 11,566.8

Therefore net gain company have in 90 days by giving discount = 11,566.8 - 4,860 = $ 6,706.8

  net gain company have in 360 days by giving discount = $ 6,706.8 x 4 = $ 26,827.2.